Find the pair of parallel lines.
(1) 12y + 6x=2
(2) 6x-12y=2
(3) -4y=-2x+2
which pair is it?
(1) and (3)
(1) and (2)
(2) and (3)
There are no parallel lines.

Answer by Alan3354(30983)

Find the pair of parallel lines.
(1) 12y + 6x=2
m = -1/2
(2) 6x-12y=2
m = +1/2
(3) -4y=-2x+2
m = +1/2
which pair is it?
(1) and (3)
(1) and (2)
(2) and (3)
2 and 3 have the same slope, m = +1/2, so they're parallel.